OpenStudy (anonymous):

Find Derivative of .... (x^2-5x+9)^.5

OpenStudy (anonymous):

\[(x^2-5x+9)^{1/2}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Im pretty positive the answer is 1/2(x2−5x+9)^1/2 x (2x-5)

OpenStudy (anonymous):

but i want to simplify it further in order to find the critical points of the derivative so I just need help with the algebra part to simplify it down...

OpenStudy (anonymous):

\[\frac{2x-5}{2\sqrt{x^2-5x+9}}\]

OpenStudy (anonymous):

zero at \[x=\frac{5}{2}\] and critical points where the denominator = 0 (which is where your original function is 0 as well.

OpenStudy (anonymous):

actually your function is never zero, so forget that. just \[x=\frac{5}{2}\]

OpenStudy (anonymous):

remember the negative sign on your exponent. \[1/2\left( x ^{2} - 5x +9\right)^{-1/2} \times \left( 2x-5\right)\]

OpenStudy (anonymous):

\[U ^{1/2}\] \[1/2U ^{-1/2}\] \[1/2(x^2-5x+9)^{-1/2}(2x-5)\]

OpenStudy (anonymous):

pretty much rewrite and solve for x

OpenStudy (anonymous):

\[x ^{1/2} =\sqrt{x}\] \[x ^{-1} = 1/x\] \[x^{-1/2} = 1/\sqrt{x}\]

OpenStudy (anonymous):

to the original poster, that's some help w/ simplifying